# Graph Representation of Narrative Context: Coherence Dependency via Retrospective Questions

### v1

@@ Abstract
This work introduces a novel and practical paradigm for narrative comprehension, stemming from the observation that individual passages within narratives are often cohesively related than being isolated. We therefore propose to formulate a graph upon narratives dubbed NarCo that depicts a task-agnostic coherence dependency of the entire context. Especially, edges in NarCo encompass retrospective free-form questions between two context snippets reflecting high-level coherent relations, inspired by the cognitive perception of humans who constantly reinstate relevant events from prior context. Importantly, our graph is instantiated through our designed two-stage LLM prompting, thereby without reliance on human annotations. We present three unique studies on its practical utility, examining the edge efficacy via recap identification, local context augmentation via plot retrieval, and broader applications exemplified by long document QA. Experiments suggest that our approaches leveraging NarCo yield performance boost across all three tasks.
Graph Representation of Narrative Context:
Coherence Dependency via Retrospective Questions
Liyan Xu  Jiangnan Li  Mo Yu  Jie Zhou Pattern Recognition Center, WeChat AI liyanlxu@tencent.commoyumyu@global.tencent.com
@@ 1 Introduction
Text comprehension has been advanced significantly ascribed to Large Language Models (LLMs), especially with long context window enabled via techniques such as positional scaling Xiong et al. (2023); Peng et al. (2024) and efficient attention Wang et al. (2023); Chen et al. (2024). Nevertheless, though extending context window could resolve certain long context tasks end-to-end, e.g. question answering, more fine-grained tasks that require explicit global dependency beyond local evidence still remain a challenge.
In book-level narrative understanding particularly, such as retrieving relevant plot passages of queries Xu et al. (2023b), or identifying recap passages of a given plot Li et al. (2024), each local passage in a novel rather serves specific purposes to other parts than being isolated, which may be easily neglected in the end-to-end process without explicit modeling these global dependency relations.
Traditionally, discourse parsing is established to capture those coherence relations between sentences, characterizing how each proposition relate to others to reflect high-level understanding of the global content Grosz and Sidner (1986). Past works have materialized various discourse frameworks, such as Rhetorical Structure Theory Mann and Thompson (1988) and Penn Discourse Treebank Prasad et al. (2008). However, despite its adoption in certain applications Bhatia et al. (2015); Ji and Smith (2017); Xu et al. (2020); Pu et al. (2023), they have not attracted appreciable focus in a wider spectrum of tasks; the underlying reasons may be twofold. First, popular discourse formalisms pose finite relation space with fixed taxonomies, offering trivial auxiliary signals especially upon LLMs. Second, they require trained experts to perform annotation for training proper parsers, hindering the overall utility due to inevitably limited resources.
In this work, we propose an alternative path to handle the aforementioned challenges in long narrative understanding, which can be deemed as a new paradigm of quasi-discourse parsing. To overcome previous limitations so to promote practical values for narrative tasks, our approach is designed to obtain flexible coherence relations without tying to formal linguistics or human annotations, thus being directly applicable as an off-the-shelf option.
@@
Drawing inspiration from the human cognitive process on narrative perception, whereas humans can constantly reinstate relevant causal events from past context during reading Trabasso and Sperry (1985); Graesser et al. (1994), our proposed idea is simple and intuitive: a NARrative COgnition graph is built, dubbed NarCo, where the entire context is split into small chunks that serve as graph nodes, and edges are connected that represent the relations between node pairs. Particularly, edges are constituted by free-form questions regarding the two connecting nodes, aligned with recent discourse works on Questions Under Discussion Kuppevelt (1995) such as DCQA Ko et al. (2022, 2023). As humans could relate to past context in retrospect, accordingly, each question in NarCo arises from the succeeding node, asking necessary background or causes that can be clarified by the preceding node. Hence, graph edges consist of inquisitive questions that naturally reflect retrospection. Overall, the resulting graph explicitly depicts task-agnostic understanding of fine-grained coherence flow that could be flexibly utilized by downstream tasks.
The key difficulty of the graph lies in the edge realization, which itself requires capable context understanding, to determine which aspects to inquire upon the context and distinguish whether they are salient for the comprehension. Such process is especially strenuous due to the large hypothesis space compared to conventional discourse formalisms, which may only become feasible recently with assistance by LLMs. To this end, we construct edges automatically through our proposed LLM prompting scheme without human annotation constraints, of which consists a question generation stage and a back verification stage (Section 3).
NarCo primarily addresses challenges for narratives on two perspectives. First, the graph edges provide a view of explicit information flow, enabling task-specific guidance towards the narrative development. Second, each chunk is now enriched with dependency of global coherence relations, thus provided augmentation of local context to deepen the digest of independent passages.
@@
To empirically demonstrate the practical utility of NarCo, we present studies on the following narrative understanding tasks from three angles.
@@
$\bullet$ Our first study examines the edge efficacy on whether it expresses competent coherence relations (Section 4). We conduct experiments on recap passage identification Li et al. (2024), where NarCo boosts up to 4.7 F1 over the GPT-4 baseline.
@@
$\bullet$ Our second study concerns the exploitation of enriched local embeddings, by injecting edges with global dependencies into node representation (Section 5). Evaluated on the plot retrieval task Xu et al. (2023b), our proposed approach with NarCo outperforms the zero-shot baseline by 3% and the supervised baseline by 2.2%.
@@
$\bullet$ Lastly, we utilize NarCo in long document question answering as a broader application (Section 6). Experiments on QuALITY that requires global evidence Pang et al. (2022) suggest that, NarCo consistently raises zero-shot accuracy by 2-5% upon retrieval-based baselines with various LLMs, able to recognize more relevant context.
Overall, our contributions can be listed as follows:
  • A new paradigm for narrative understanding is proposed, parsing the context into a graph of high-level coherence relations, named NarCo.
  • The graph is practically realized with our two-stage LLM prompting w/o human annotations.
@@
  • We present three studies effectively utilizing NarCo on narratives, focusing on edge efficacy, node augmentation, and broader applications.
@@ 2 Related Work
QUD is a linguistic framework with rich history that approaches discourse and pragmatics analysis by repeatedly resolving queries triggered by prior context Kuppevelt (1995); Roberts (1996); Benz and Jasinskaja (2017). Recent works have begun adapting QUD for discourse coherence De Kuthy et al. (2018, 2020); Ko et al. (2020, 2022, 2023) or other applications Wu et al. (2023b); Newman et al. (2023). Our proposed NarCo can also be perceived as a unique form of QUD variant, though it is principally rooted upon narrative comprehension rather than formal linguistics. Therefore, NarCo differs from QUD works considerably on the following design choices.
Coarse Granularity   While QUD tends to employ sentences as the basic discourse unit, NarCo opts for a coarser granularity, adopting passages (or short chunks) as graph nodes. It is driven by the fact that in narratives, complex events or interactions may often be conveyed beyond sentence-level, thus relations in NarCo could target higher-level understanding between nodes.
Retrospection-Oriented   Unlike conventional QUD that inquires from prior context to be addressed by subsequent context (forward direction), which could yield unanswerable questions Westera et al. (2020); Ko et al. (2020), NarCo takes the backward direction, by asking retrospective questions from succeeding context, such that all generated questions in NarCo are naturally grounded by corresponding prior context.
Precision-Focused   Unlike previous QUD works that require dedicated human annotations, NarCo is formulated attainable by LLMs. Accordingly, we prioritize precision over recall for practical instantiation of graph edges, and do not necessitate strict linguistic criteria, as long as it contributes positively for narrative understanding.
@@
A major task direction on narratives is question answering (QA), where past works have proposed several datasets with human annotations, such as NarrativeQA Kočiský et al. (2018), TellMeWhy Lal et al. (2021), FairytaleQA Xu et al. (2022b), QuALITY Pang et al. (2022). We choose QuALITY as the broader application in this work, due to its challenging long context, requirement of global evidences, and simple evaluation by multi-choices.
Recently, several tasks have emerged focusing on modeling the reading process of long narratives, including TVShowGuess Sang et al. (2022), PersoNet Yu et al. (2023), ToM-in-AMC Yu et al. (2024) and retrieval tasks studied in this work such as RELiC Thai et al. (2022), PlotRetrieval Xu et al. (2023b). These tasks require a holistic understanding of the long narratives to enhance contextual comprehension of specific segments. We recognize the significance of explicitly modeling global coherence dependency as a crucial aspect of narrative comprehension, thereby motivating the inception of our research.
Long Context Understanding
One of the major research directions of LLMs is the extension of context window, which can be seamlessly applied for long context understanding tasks, including scaling positional embeddings Chen et al. (2023b); Xiong et al. (2023); Peng et al. (2024), efficient attention Zaheer et al. (2020); Chen et al. (2024), cached attention Wu et al. (2022); Wang et al. (2023), recurrent attention Dai et al. (2019), etc. Though effective, many narrative tasks demand beyond the end-to-end solution. Recently, new paradigms have been proposed for fine-grained processing, such as compressing context as soft prompts Chevalier et al. (2023), and MemWalker that reads long context interactively via iterative prompting Chen et al. (2023a). Nevertheless, our proposed approach takes parsing as an alternative paradigm, which is orthogonal to the existing directions and could be even further combined in the future.
@@
Utilizing various structures in natural languages has attracted much attention by previous works, such as syntactic structures on token-level Strubell et al. (2018); Xu et al. (2022a), relation structures on span-level Xu and Choi (2020, 2022); Nguyen et al. (2022) or discourse structures on span or sentence-level Bhatia et al. (2015); Ji and Smith (2017). In addition, there have been studies constructing graphs with entity nodes Min et al. (2019); Ding et al. (2019) based on Wikipedia links for factoid question answering. Li et al. (2020, 2021) have extended these graphs by incorporating event nodes. As all these structures encompass pre-defined taxonomies on edge types, our propose graph representation is motivated to comprise open-world edge types that have been practiced in other tasks Wu et al. (2019); Xu et al. (2023a); Su et al. (2024), meanwhile attainable by LLMs without efforts of human annotations.
@@ 3 NarCo: Narrative Cognition Graph
In this section, we start by delineating our graph formulation, which is itself not tied to any particular implementation. Subsequently, we elaborate our graph realization using LLMs, as our endeavor to remove dependence on human annotations.
@@ 3.1 Graph Formulation
An edge connecting two nodes signifies the relationships they convey. These relations are articulated as free-form inquisitive questions that are not constrained by fixed taxonomies. All edges follow the backward direction, such that for an edge $e_{ij}$ ($i<j$), the expressed questions always arise from the succeeding node $v_{j}$, asking clarification regarding specific events or situations appeared in $v_{j}$, which could be addressed by the preceding $v_{i}$. For narratives, questions primarily target causal and temporal relations as the coherence dependency.
Functionally speaking, these backward edges resemble the human cognitive process for narrative perception: when reading a certain passage, humans are able to reinstate previous relevant parts in retrospect that lay out the build-up or causes, so to achieve a causally coherent comprehension of the global context Trabasso and Sperry (1985); Graesser et al. (1994); Song et al. (2020). Unlike QUD that originally features curiosity-driven questions in a forward direction, which could yield unanswerable questions, all edges in NarCo are fully grounded by the context, such that all questions are addressable by prior nodes, thereby serving as the bridge for global coherent dependency.
@@
  • It may have zero-to-many questions. An empty edge without questions indicates the succeeding node $v_{j}$ is independent from $v_{i}$ in terms of causal or temporal relations.
  • Each question should be salient towards the comprehension of narrative development, rather than being trivial details. Hence, the number of questions of an edge should reflect how coherently related between the two nodes.
  • As we adopt coarse granularity for nodes, questions could inquire higher-level relations based on the extrapolation over multiple sentences, which may be useful for downstream tasks.
@@ 3.2 Graph Realization
To obtain graph nodes, the full context is split by paragraphs and sentences. We impose each node within 240 words in this work, though the exact limit can be task-specific. For a graph characterized by $N$ total nodes, there are $O(N^{2})$ edges available, which can become cumbersome and excessive. It is also task-dependent to determine which pairs of nodes should be gathered edges upon, e.g. for enriching local representation, it may be enough to obtain proximal coherence relations by edges from neighboring nodes within a context window.
Despite the daunting task of edge construction, the emergence of LLMs presents an opportunity: through LLM prompting, it becomes conceivable to actualize the entire graph without any human annotations involved. To this end, we introduce a two-stage prompting scheme as follows for the challenging edge formulation in NarCo.
@@
For an edge $e_{ij}$ to be instantiated, LLM needs to determine important aspects to ask upon $v_{j}$ that reflect the retrospective coherence towards the prior context in $v_{i}$. Similar utilization of LLMs for question generation (QG) has been explored in other applications, such as performing QG for QUD Wu et al. (2023a) and passage decontextualization Newman et al. (2023), where LLM is prompted to generate questions directly based on task-specific criteria. For our case, such direct generation can be briefly outlined as:
@@
However, our preliminary experiments found that, though LLM could follow the instructions to generate plausible questions, their quality is often unsatisfactory for NarCo, with common errors as follows (examples in Appx A.2):
@@
  • LLM often asks questions upon $v_{j}$ but also answerable by $v_{j}$ as well. Such patterns align with the more conventional QG setting Du et al. (2017) that may exist plentifully during supervised finetuning of LLM. However, they are not desirable for NarCo as they cannot effectively indicate coherence with $v_{i}$.
@@
  • LLM could hallucinate their relations by guessing and inferring extra underlying connections not grounded by the provided context, resulting in questions not directly answerable by $v_{i}$.
@@
In essence, QG for NarCo requires LLM simultaneously aware of questions being: 1) arising from $v_{j}$; 2) not answerable by $v_{j}$; 3) answerable by $v_{i}$. As this process is empirically challenging even for strong LLM (e.g. GPT-4), we perform QG with two heuristic turns that could be viewed as Chain-of-Thoughts Wei et al. (2022) guided by prompts:
  1. 1.
@@
    List concrete parts in $v_{i}$ that contribute as the preceding background or cause for specific events or situations mentioned in $v_{j}$, along with brief explanations.
  2. 2.
@@
    Convert each listed connection to a question, such that it asks about the cause or background upon $v_{j}$ and can be clarified by the corresponding concrete part in $v_{i}$, helpful to comprehend their causal or temporal relations.
Question Filtering
@@
Our pilot study suggests that the above two-turn QG could yield higher-quality questions than rudimentary generation, especially reducing self-answerable questions. However, it is still inevitable to produce noisy questions of the two identified error types. In light of remaining noises, we propose a second stage to filter noisy questions through back verification, akin to the concept of back translation Sennrich et al. (2016):
@@
Given a context $\mathcal{C}_{ij}$ and a related question, determine whether it can be answered. If so, reason the answer and provide original sentences of key supporting evidence.
@@
Particularly, $\mathcal{C}_{ij}$ is the concatenated context from $v_{i}$ and $v_{j}$ without disclosing their boundary. If the question is answerable, we then parse the response and identify whether the supporting sentences are from $v_{i}$. If not, the question becomes invalid and discarded, as it does not offer to bridge two nodes.
@@
Overall, all generated questions are back verified; only questions that could be answered by prior nodes are finally retained in NarCo, being a precision-focused approach. In this work, we adopt GPT-4 for strong question generation, and ChatGPT for the easier verification. NarCo may also be obtained by strong open-source LLM or trained models distilled from GPT-4. Our full prompts and more details are provided in Appx A.1.
3.3 Graph Analysis
Upon examination of preliminary results on the English version of the novel Notre-Dame de Paris, edges in NarCo mostly encompass what and why types of questions, approximately constituting 61% and 26%. It is worth noting that many questions reflect high-level understanding of the context (examples in Appx A.2), in contrast to conventional discourse relations, e.g. purpose, condition in RST Mann and Thompson (1988). With the two-stage prompting scheme, the averaged node degree reaches 1.9. Particularly, the verification stage identifies 47.4% questions to be filtered out.
As our graph formulation primarily aims at practical values, we demonstrate its effective utility in Sections 4-6, addressing three distinct perspectives.

### v2

# Fine-Grained Modeling of Narrative Context: A Coherence Perspective via Retrospective Questions

@@ Abstract
This work introduces an original and practical paradigm for narrative comprehension, stemming from the characteristics that individual passages within narratives tend to be more cohesively related than isolated. Complementary to the common end-to-end paradigm, we propose a fine-grained modeling of narrative context, by formulating a graph dubbed NarCo, which explicitly depicts task-agnostic coherence dependencies that are ready to be consumed by various downstream tasks. In particular, edges in NarCo encompass free-form retrospective questions between context snippets, inspired by human cognitive perception that constantly reinstates relevant events from prior context. Importantly, our graph formalism is practically instantiated by LLMs without human annotations, through our designed two-stage prompting scheme. To examine the graph properties and its utility, we conduct three studies in narratives, each from a unique angle: edge relation efficacy, local context enrichment, and broader application in QA. All tasks could benefit from the explicit coherence captured by NarCo.
[…]
[ leftline=true, rightline=false, topline=false, bottomline=false, linecolor=gray, linewidth=2pt, innerrightmargin=7pt, innerleftmargin=8, innertopmargin=3pt, innerbottommargin=3pt, leftmargin=6pt, rightmargin=5pt, skipabove=skipbelow=]leftbarquote
Fine-Grained Modeling of Narrative Context:
A Coherence Perspective via Retrospective Questions
Liyan Xu   Jiangnan Li   Mo Yuthanks: Corresponding author.   Jie Zhou Pattern Recognition Center, WeChat AI {liyanlxu,jiangnanli,withtomzhou}@tencent.commoyumyu@global.tencent.com
@@ 1 Introduction
Since the advent of Large Language Models (LLMs), document comprehension has been improved significantly by simply employing the end-to-end generative paradigm. Especially, with long context window enabled via techniques such as position interpolation Xiong et al. (2023); Peng et al. (2024), cached or efficient attention Wang et al. (2023); Ge et al. (2024a); Munkhdalai et al. (2024), context compression or pruning Chevalier et al. (2023); Anagnostidis et al. (2023); Ge et al. (2024b), the end-to-end paradigm is deemed undoubtedly simple and effective for comprehension tasks (e.g. question answering) on various documents.
However, while the typical benchmarks are continually enhanced by more advanced LLMs Touvron et al. (2023); Jiang et al. (2024); OpenAI et al. (2024), the end-to-end paradigm may not suffice for all comprehension scenarios. In this work, we focus around the narrative context, i.e. stories or novels, and propose a conceptually original framework of fine-grained context modeling: a graph is formulated that depicts the relations between context snippets, abstracting over the context to reflect a high-level understanding of the narrative. The graph itself is practically realized by LLMs to harness their rapidly evolving strengths, and the resulting graph could serve to facilitate various downstream narrative comprehension tasks.
Our motivation arises from the distinctive nature of narratives: multiple development of characters or events in a story could be entangled over long context ranges, where each local passage usually serves specific purposes for others. Thus, individual passages tend to be cohesively interconnected than being isolated. As the end-to-end paradigm implicitly grasps these context connections through sequence modeling, our approach explicitly models these dependency relations to capture coherence, offering a directly-applicable alternative path orthogonal to the end-to-end paradigm.
@@
Concretely, drawing inspiration from the cognitive process on narrative perception, whereas humans can constantly reinstate relevant or causal events from past context during reading Trabasso and Sperry (1985); Graesser et al. (1994), our formalism, termed NARrative COgnition graph (NarCo), splits the entire context into chunks that act as graph nodes, with edges representing the relations between node pairs. In particular, edge relations are constituted by free-form questions. As humans could relate to past context in retrospect, accordingly, each question in NarCo edges arises from the succeeding node (latter context), asking necessary background or causes that can be clarified by the preceding node (prior context). Hence, graph edges consist of inquisitive questions that naturally reflect retrospection. Overall, the resulting graph explicitly depicts task-agnostic understanding of fine-grained coherence flow that could be flexibly utilized by downstream tasks.
Though our graph formulation partially shares motivations with discourse parsing that characterizes how each proposition relates to others within a close context Grosz and Sidner (1986), our method targets on practical utility for narrative comprehension, where edges in NarCo are designed to be easily obtained and effectively consumed by downstream tasks. Consequently, NarCo is formulated in a different scope from discourse parsing by two main perspectives. First, as most discourse frameworks, such as Rhetorical Structure Theory Mann and Thompson (1988), Penn Discourse Treebank Prasad et al. (2008), or the recent Questions Under Discussion (QUD) Ko et al. (2022, 2023) are rooted upon linguistic principles, their relation types are oriented for formal discourse analysis, requiring trained experts to annotate edges according to a defined linguistic taxonomy. Whilst for NarCo, the relation space is larger without taxonomy constraints, offering diverse high-level semantic signals for narrative tasks. Second, NarCo practically leverages LLMs to derive edge relations, without reliance on human annotations. Thus, the edge quality is not restricted by annotation resources, and shall be continuously enhanced along with the ongoing LLM advancement.
The key difficulty of NarCo lies in the edge realization between two nodes, which itself demands strong context understanding to determine which aspects to inquire upon the context, and to assess their saliency for comprehension. Such process is especially strenuous due to the large hypothesis space compared to conventional discourse formalisms. To this end, we pose soft semantic constraints on relations, and employ LLM’s capabilities to construct edges automatically through our proposed prompting scheme, of which consists a question generation stage and a self verification stage (Section 3). The obtained edges could then be utilized by downstream tasks in two primary ways. First, edges themselves directly provide information flow to guide various comprehension tasks. Second, they offer global coherence view for each node, thereby augmenting the local context to deepen the digest of independent passages.
@@
To empirically demonstrate the practical utility of NarCo, we present three studies on narrative comprehension tasks, each from a unique angle:
@@
$\bullet$ Our first study examines the edge efficacy on whether the relation questions capture capable retrospective coherence (Section 4). We conduct experiments on the recap identification task Li et al. (2024), where NarCo is shown to recognize coherence dependencies between context, boosting up to 4.7 F1 over the GPT-4 baseline.
@@
$\bullet$ Our second study concerns the exploitation of enriched local embeddings, by injecting edges of relation dependencies into node representation (Section 5). Evaluated on the plot retrieval task Xu et al. (2023b), our proposed approach with NarCo outperforms the zero-shot baseline by 3% and the supervised baseline by 2.2%.
@@
$\bullet$ Lastly, we utilize NarCo in a long document question answering task (Section 6), as a broader application of Retrieval-Augmented Generation (RAG) Lewis et al. (2020). Experiments on QuALITY that requires global context evidence Pang et al. (2022) suggest that, NarCo consistently raises zero-shot accuracy by 2-5% upon retrieval-based baselines with various LLMs, able to identify more relevant context through edge relations.
Overall, our key contributions in this work are:
  • We propose a new paradigm of fine-grained context modeling to facilitate narrative comprehension, orthogonal to the end-to-end paradigm.
  • Our introduced NarCo framework describes flexible relations of context dependencies by retrospective questions, which are realized by LLMs through our designed prompting scheme, without reliance on human annotations.
@@
  • We present three studies effectively utilizing NarCo on narrative tasks, empirically examining its edge properties and broader utilization.
@@ 2 Related Work
QUD is a linguistic framework with rich history that approaches discourse and pragmatics analysis by repeatedly resolving queries triggered by prior context Kuppevelt (1995); Roberts (1996); Benz and Jasinskaja (2017). QUD has been adapted by recent works for discourse analysis De Kuthy et al. (2018, 2020); Ko et al. (2020, 2022, 2023) or other applications Wu et al. (2023b); Newman et al. (2023). Our proposed NarCo also adopts question-form relations; though, the scope and motivation is different from discourse analysis. Consequently, NarCo differs from QUD works considerably on the following design choices.
Coarse Granularity   While QUD tends to employ sentences as the basic discourse unit, NarCo opts for a coarser granularity, adopting passages (or chunks) as the graph nodes. It is driven by the fact that in narratives, complex events or interactions may often be conveyed beyond sentence-level, thus relations in NarCo could target higher-level understanding between context.
Retrospection-Oriented   Unlike conventional QUD that inquires from prior context to be addressed by subsequent context (forward direction), which could yield unanswerable questions Westera et al. (2020); Ko et al. (2020), NarCo takes the backward direction, by asking retrospective questions from latter context, such that all generated questions in NarCo are naturally grounded by the corresponding prior context.
Precision-Focused   Unlike previous QUD works that require dedicated human annotations, NarCo is formulated attainable by LLMs. Accordingly, we prioritize precision over recall for practical instantiation of graph edges, and do not necessitate strict linguistic criteria, as long as edges contribute positively for narrative understanding.
@@
A major task direction on narratives is question answering (QA), where past works have proposed several datasets with human annotations, such as NarrativeQA Kočiský et al. (2018), TellMeWhy Lal et al. (2021), FairytaleQA Xu et al. (2022b), QuALITY Pang et al. (2022). We adopt QuALITY as the broader application in this work, due to its challenging long context, requirement of global evidences, and simple evaluation by multi-choices.
Recently, several tasks have emerged focusing on modeling the reading process of long narratives, including TVShowGuess Sang et al. (2022), PersoNet Yu et al. (2023), ToM-in-AMC Yu et al. (2024), and retrieval tasks such as RELiC Thai et al. (2022), PlotRetrieval Xu et al. (2023b). These tasks require a holistic understanding of the long narratives to enhance contextual comprehension of specific segments. We reckon the significance of explicitly modeling context dependencies as a crucial aspect of narrative comprehension, motivating the inception of .
LLM Understanding and Reasoning
LLMs have demonstrated remarkable capabilities on a wide spectrum of comprehension and reasoning tasks Chen et al. (2024); Sun et al. (2024). The simple end-to-end solution is especially appealing with long context window enabled, using techniques such as scaling positional embeddings Chen et al. (2023b); Xiong et al. (2023); Peng et al. (2024), efficient attention Munkhdalai et al. (2024), cached attention Wang et al. (2023); Ge et al. (2024a), recurrent attention Dai et al. (2019), context compression Chevalier et al. (2023); Ge et al. (2024b), context pruning Anagnostidis et al. (2023), etc. Though being effective, certain narrative tasks demand beyond the end-to-end solution. Recently, new methods have been proposed for fine-grained task processing, e.g. reading agents such as MemWalker Chen et al. (2023a). Nevertheless, our proposed approach depicts explicit context dependencies as an alternative paradigm, which is orthogonal to the existing directions and could be even further combined.
@@
Various relational structures in text documents has attracted much attention by previous works, such as syntactic relations Strubell et al. (2018); Xu et al. (2022a), discourse relations Ji and Smith (2017); Nair et al. (2023); Hu and Wan (2023), entity or event relations Ding et al. (2019); Li et al. (2020, 2021); Xu and Choi (2020, 2022); Nguyen et al. (2022). As all these structures encompass pre-defined taxonomies on edge types, our propose graph representation is motivated to comprise open-world edge types that have been practiced in other tasks Wu et al. (2019); Xu et al. (2023a); Su et al. (2024), while being practical and attainable by LLMs without requiring efforts of human annotations.
@@ 3 NarCo: Narrative Cognition Graph
In this section, we start by delineating our graph formulation, which is itself not tied to any particular implementation. Subsequently, we elaborate our specific graph realization using LLMs, without dependence on human annotations.
@@ 3.1 Graph Formulation
An edge connecting two nodes indicates the relations between the context. These relations are articulated as free-form questions that are not constrained by fixed taxonomies. All edges follow the backward direction, such that for an edge $e_{ij}$ ($i<j$), the expressed questions always arise from the succeeding node $v_{j}$, asking clarification regarding specific events or situations appeared in $v_{j}$, which could be addressed by the preceding context $v_{i}$. Since the hypothesis space is huge without any regularization, we pose soft semantic constraints on questions, such that questions should primarily reflect on causal and temporal relations, which are significant to the narrative coherence.
Functionally speaking, these backward edges resemble the human cognitive process for narrative perception: when reading a certain passage, humans are able to reinstate previous relevant parts in retrospect that lay out the build-up or causes, so to achieve a coherent comprehension of the global context Trabasso and Sperry (1985); Graesser et al. (1994); Song et al. (2020). Unlike conventional QUD that features curiosity-driven questions in a forward direction, which could yield unanswerable questions, all edge in NarCo are fully grounded by the prior context, since all retrospective questions are addressable by prior nodes.
@@
  • It may have zero or many questions. An empty edge (zero questions) signifies $v_{j}$ is vaguely independent from $v_{i}$ in terms of coherence.
  • Each question should be salient towards the comprehension of narrative development, rather than inquiring trivial details. Hence, the number of questions in an edge should reflect how cohesively related between two nodes.
  • As we adopt coarse granularity for nodes, questions could probe higher-level relations based on the extrapolation over multiple sentences, which may be useful towards broader understanding.
@@ 3.2 Graph Realization
To obtain graph nodes, the full context is split by paragraph and sentence boundaries. We impose each node within 240 words in this work, though the exact limit can be task-specific. For a graph characterized by $N$ total nodes, there are $O(N^{2})$ full edges available, which can become cumbersome and excessive. It is also task-dependent to determine which pairs of nodes should be gathered edges upon, e.g. for enriching local representation, it is sufficient to obtain relation dependencies from neighboring nodes within a context window.
Despite the daunting formulation on edge relations, the emergence of LLMs presents an opportunity: through designed LLM prompting, it becomes conceivable to actualize the entire graph without any human annotations involved. To this end, we introduce a two-stage prompting scheme to tackle the challenging edge construction.
@@
For an edge $e_{ij}$ to be instantiated, LLMs need to determine important aspects to ask upon $v_{j}$ that reflect the retrospective coherence towards the prior context in $v_{i}$. Similar utilization of LLMs for question generation (QG) has been explored in other applications, such as performing QG for QUD Wu et al. (2023a) and passage decontextualization Newman et al. (2023), where a LLM is prompted to generate questions directly based on task-specific criteria. For our case, such direct generation can be briefly outlined as:
{leftbarquote}
@@
However, our preliminary experiments suggest that although LLMs can generate plausible questions by following the instructions, their quality is often unsatisfactory for NarCo, with common errors as follows (examples in Appx A.2):
@@
  • Self-answerable: LLMs often ask questions upon $v_{j}$ but also answerable by $v_{j}$ as well. Such self-answerable pattern aligns with the more conventional QG setting Du et al. (2017) that may exist plentifully during the supervised finetuning of LLMs, causing a bias towards this type of question. However, they are not desirable for NarCo, since they do not express dependencies between nodes to reflect their relations.
@@
  • Hallucination: LLMs could hallucinate the relations of two nodes by guessing and inferring extra underlying connections not grounded by the provided context, resulting in questions not directly answerable by $v_{i}$.
@@
In essence, QG for NarCo requires LLM simultaneously aware of questions being: 1) arising from $v_{j}$; 2) not answerable by $v_{j}$; 3) answerable by $v_{i}$. As this is empirically challenging even for strong LLMs (e.g. GPT-4), we perform QG with two heuristic turns that can be viewed as human-guided Chain-of-Thoughts Wei et al. (2022):
{leftbarquote}
@@
1.  List concrete parts in $v_{i}$ that contribute as the preceding background or cause for specific events or situations mentioned in $v_{j}$, along with brief explanations.
@@
2.  Convert each above listed connection to a question, such that it asks about the cause or background upon $v_{j}$ and can be clarified by the corresponding concrete part in $v_{i}$, helpful to comprehend their causal or temporal relations.
@@
The designed two-turn QG scheme yields higher-quality questions than the rudimentary generation, mainly alleviating the self-answerable problem. However, noisy questions of the two identified error types still occur due to imperfect instruction following by LLMs. In light of these noises, we apply an optional second stage to filter out noisy questions through self verification.
Self Verification
The second stage takes the generated questions from QG and in turn, performs question answering on the context:
{leftbarquote}
@@
Given a context $\mathcal{C}_{ij}$ and a related question, determine whether it is answerable. If yes, reason the answer and provide original sentences of key supporting evidences.
@@
In Particular, $\mathcal{C}_{ij}$ is the concatenated context from $v_{i}$ and $v_{j}$ without disclosing their boundary. If the question is answerable, we then parse the response and identify whether the supporting sentences are from the prior context $v_{i}$. If not, the question is attested noisy and gets discarded, as it does not bridge the two context effectively.
@@
With the second stage, only questions that could be answered by prior nodes are eventually retained in NarCo, being a precision-focused approach. In this work, we adopt GPT-4 for the challenging QG stage, and ChatGPT for the easier verification stage. NarCo may also be derived with strong open-source LLMs as well. Our full prompts and more details are provided in Appx A.1.
As NarCo targets the practical utility to facilitate narrative comprehension, the obtained edges shall be directly consumed by downstream tasks. Sections 4-6 present three empirical studies, each from a distinctive perspective, to examine the edge properties and their utilization.